OpenStudy (anonymous):

3. Given that one of the root is 3 times the others for the quadratic equation 3x^2-2x+p=0, find a)the value of p b)the two roots pls check wether my working and answer is correct @ganeshie8

OpenStudy (anonymous):

use general formula of quadratic equation ?

OpenStudy (anonymous):

\[\frac{ 2 }{ 3 }\div \frac{ 4 }{ 3 }=\frac{ 1 }{ 2 }\]

OpenStudy (mrnood):

Using the quadratic equation will give you 2 equation (using the +/-) each linear eqations in p solution 1 = 3xsolution 2 so write that equation which will be a simple equation in p solve for p then solve the 2 roots from the original quadratic formula

OpenStudy (anonymous):